• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 00:01:37 on June28,2005

    Entry number 147298

    keyword=Shift summary

    Crew: E.Chudakov, N.Lianage, P.Pamela
    16:00 Status: control access to fix vacuum problems in the chicane area.
    18:25 We asked for beam permit.
    19:10 CW beam in the hall (C target, 5µA)
    19:20 55 µA, production start
    20:10 MCC: the Moller quads are off. Started to investigate.
    20:40 We resumed production with the quads off (it seems OK)
    21:40 Beam is taken away to tune Hall C
    23:25 Beam is back


    run target comment
    4921 C-thin short run to check the raster
    4922 H2O 1M wrong prescaling factors (a lot of T1...)
    4223 same 0.4M MCC noticed that the Moller quads were OFF! run stopped.
    4926 same 1M production (Moller quads off)
    4927 same 0.08M - run stopped since MCC took the beam away
    4928 same - running -